OpenStudy (anonymous):

The ratio of the surface area of two spheres is 3:2. The volume of the larger sphere is 2,916 in^3. What is the volume of the smaller sphere?

OpenStudy (saifoo.khan):

The equation for the surface area of a sphere is A = 4 pi r^2. The equation for the volume of a sphere is V = 4/3 pi R^3. Say R is the radius of the bigger sphere. Then it has surface area: 4 pi R^2 It has volume: 4/3 pi R^3 Say r is the radius of the smaller sphere. Then it has surface area: 4 pi r^2 It has volume: 4/3 pi r^3 The ratio between the surface areas is then just the surface area of the bigger one divided by the surface area of the smaller one: (4 pi R^2) / (4 pi r^2) = R^2 / r^2 The question says that this is 3:2, which means that: R^2 / r^2 = 3 / 2. We can get rid of the fractions by multiplying diagonally through the equation: 3r^2 = 2R^2 Now the volume of the larger sphere is 2.916. We can plug this into the volume equation to get the radius of the larger sphere, which is R: V = 4/3 pi R^3 2.916 = 4/3 pi R^3 R^3 = (3 * 2.916) / (4 * pi) R^3 = 0.69614 R = 0.8863 inches Now we can plug this into our surface area ratio equation to get r, the radius of the smaller sphere: 3r^2 = 2R^2 Now substitute in R: 3r^2 = 2 * (0.8863)^2 3r^2 = 1.57095 r^2 = 0.52365 r = 0.72364 inches Now we can use this to calculate the volume for the smaller sphere using the sphere volume equation: V = 4/3 pi r^3 V = 4/3 pi (0.72364)^3 V = 1.58727 inches^3

OpenStudy (anonymous):

5